OpenStudy (anonymous):

1162/x = 3424/4000 X=?

OpenStudy (anonymous):

Cross multiply

OpenStudy (anonymous):

1162 / x = 0.856 x = 1162 / 0.856 = ??????

OpenStudy (anonymous):

3424x=4000*1162 So multiply the 4000 by 1162 and then divide both sides by 3424